You want to determine the opinion of students in New York University about the new platform used in learning this pandemic. You wanted to use stratified sampling in getting responses from a sample of 500 students. Given the following data, identify how many samples are there in each grade level.


Grade 7: 1,557


Grade 8: 1, 209


Grade 9: 1,054


Grade 10: 1, 988


Grade 11:1, 456


Grade 12: 1,764


1
Expert's answer
2022-04-11T17:24:32-0400

There are 1557+1209+1054+1988+1456+1764=9028 students.

Samples in each grade:

Grade 7: "n=500*\frac{1557}{9028}=86."

Grade 8: "n=500*\frac{1209}{9028}=67."

Grade 9: "n=500*\frac{1054}{9028}=58."

Grade 10: "n=500*\frac{1988}{9028}=110."

Grade 11: "n=500*\frac{1456}{9028}=81."

Grade 12: "n=500*\frac{1764}{9028}=98."